# EE-LLM: Large-Scale Training and Inference of Early-Exit Large Language Models with 3D Parallelism

### 08 Dec 2023 v1

@@ 7.2 Inference
In the experiments below, we verify the effectiveness of our pipeline-based approach (with the number of pipeline stages set to 4) proposed in Section 5.2, and the method of KV recomputation explained in Section 6.3, for autoregressive generation with early exiting. Given that both methods generate the same output for the same prompt, we first investigate the downstream performance and early-exit speedup for the pipeline-based method alone, and then compare the inference speed of both methods.
To examine the performance of inference with various sizes of models, we conduct experiments under different settings. One model that we use is an 1.3B early-exit model with tied embedding matrices, pre-trained from scratch with early-exit loss weights 1/4 and 1/2 using 300B tokens. The other one is a 7B early-exit model with untied embedding matrices, pre-trained with early-exit loss weights 0.1 and 0.2 using 150B tokens. Each of them has two minimalistic early exits added to the 1/4 and 1/2 depth respectively. Neither model has gone through any fine-tuning or alignment. We adopt greedy decoding and a simple confidence-based decision rule for early exiting: at each early exit, if the maximum probability of next-token prediction is above a pre-defined threshold, then the model chooses the most likely token and moves on to generate the next one. The trade-off between output quality and speed of generation is then controlled by the threshold. Note that for a fair comparison, when the threshold is set to 1, early-exit layers are disabled, so that the computational complexity of a full-model forward pass matches that of a standard LLM without early exits; the inference latency with the threshold set to 1 is also used as the baseline for calculating (relative) speedup.
@@ 7.2.1 Output quality and speed of generation
We evaluate the downstream performance of our models with HELM [22] in four tasks. Among them, BoolQ [8] and TruthfulQA [23] are question-answering tasks, while XSUM [25] and CNN/DailyMail [39] are summarization tasks. We use EM (ratio of exact match with the correct answers) as the metric for the former, and ROUGE-L (a measure of similarity to the reference summaries) for the latter. Figure 9 demonstrates the relation between scores and inference speedup (with varying thresholds) for pipeline-based inference. Encouragingly, for many of these tasks, inference with early exiting can achieve 2×2$\times$2 × or higher speedup compared to full-model inference, with comparable or even better evaluation scores. Figure 10 provides a concrete piece of text generated by our 7B early-exit model, showing that speedup can be achieved with minor or no impact on the resulted sequence. Figure 11 goes one step further and demonstrates the confidence at each exit for each token, confirming the existence of easy tokens that can be predicted with high confidence at early exits, without wasting computational resources on later layers. We conjecture that early exiting can, in principle, bring higher speedup for larger models, as expressivity of early exits grows with the model capacity, which allows them to make confident and accurate predictions for a larger proportion of tokens.
@@
Figure 9: Evaluation scores and speedup with our early-exit models. For each curve, the confidence threshold decreases from left to right.
Figure 10: Sequences generated by our 7B early-exit model for the same prompt, with varying confidence threshold. Differences are highlighted in red.
Figure 11: Confidence of next-token prediction at each exit of our 7B early-exit model for generating each token within one sequence. Tokens that are predicted with confidence higher than 0.8 are highlighted; interestingly, all exits make the same prediction for each of these tokens.
@@ 7.2.2 A comparison between two inference methods
We compare the inference time of both pipeline-based method and KV recomputation. Since the pipeline-based method uses 4 GPUs (for 4 pipeline stages), we allow KV recomputation to use a tensor parallelism (TP) degree of 1 or 4 for a fair comparison. The empirical results on XSUM and CNN/DailyMail tasks are shown in Figure 12. We first notice that the pipeline-based approach outperforms KV recomputation with $\text{TP}=1$ for confidence thresholds smaller than 1 (i.e. when early exiting actually happens), despite the point-to-point communication overhead. Moreover, KV recomputation with $\text{TP}=4$ is faster than the pipeline-based approach; it is worth noting, however, that the small speedup from $\text{TP}=1$ to $\text{TP}=4$ for KV recomputation is only possible with high-end hardware like A100 GPUs connected by high-bandwidth communication. In many use cases of LLM inference, such hardware is not available, and pipeline parallelism is the only option for partitioning a model that is too large to fit into the memory of one single device. In sum, the proposed pipeline-based approach is a more practical option for LLM inference with model partitioning.
@@ 8 Conclusions
We have introduced EE-LLM, a system for large-scale training and inference of early-exit LLMs with 3D parallelism. For training, we have presented how to execute backpropagation of the early-exit training objective across pipeline stages, various performance optimizations for minimizing the computational overhead compared to training standard LLMs, and some advanced features for more fine-grained control and optimization of the training process. For inference, we have introduced our design and implementation of two inference methods, one based on KV recomputation and the other based on a new type of pipeline parallelism, both of which are compatible with KV caching for autoregressive generation. Along the way, we have discovered many interesting connections between early-exit models and pipeline parallelism. We hope that EE-LLM will be a helpful tool for future study and applications of early-exit LLMs at larger scales.
@@
Appendix A Appendix
@@ A.1 Methodology
The maximum number of microbatches that can be inserted into Part 1 or 2 of the explicit bubbles, without increasing training time per iteration, is $\lfloor(p-1)b/(f+b)\rfloor=\lfloor(p-1)/(f/b+1)\rfloor$, where $f/b$ is (an estimate of) the ratio between forward and backward time. To see this for Part 1 (resp. 2), one simply need to notice that the first (resp. last) pipeline stage has an bubble size $b(p-1)$, while the total forward and backward time for each microbatch is $f+b$; dividing the first value by the second one concludes the proof.
@@
With $K$ inserted microbatches, the $i$-th microbatch is supposed to go through the forward pass of the first $K+1-i$ stages; however, if there is no early exit on Stage $K+1-i$, then this microbatch only need to go through the forward pass up to the last stage (among the first $K+1-i$ stages) that has at least one early exit.
@@
We can calculate the maximum number of backward stages for each inserted microbatch, while ensuring that no overhead to training time per iteration occurs. Notice that for the $i$-th microbatch, the remaining bubble size after its forward computation at the last stage is $(p-1)b-f-(i-1)(f+b)=pb-i(f+b)$; dividing this value by $b$ gives the number of backward stages $\lfloor(pb-i(f+b))/b\rfloor=\lfloor p-i(f/b+1)\rfloor$.
@@
The actual implementation of this method in EE-LLM takes a different perspective. Instead of inserting additional microbatches into the original schedule, we keep the number of microbatches per training iteration unchanged, and do the following: (1) we replace the full forward/backward computation of a few microbatches with partial forward/backward computation, which can be placed in the bubble between the warm-up and steady phases; (2) we truncate the backward computation of the last few microbatches. These two steps correspond exactly to Parts 1 and 2 introduced earlier, and the visualization in Figure 5 remains valid. Such an implementation reduces the training time per iteration, at the cost of lower data utilization.
A.2 Analysis

### 01 Feb 2024 v2

@@ 7.2 Inference
In the experiments below, we verify the effectiveness of our pipeline-based approach (with the number of pipeline stages set to 4) proposed in Section 5.2, and the method of KV recomputation explained in Section 6.3, for autoregressive generation with early exiting. Given that both methods generate the same output for the same prompt, we first investigate the downstream performance and early-exit speedup for the pipeline-based method alone, and then compare the inference speed of both methods. A server with 4 Nvidia A100-40GB GPUs is used for inference.
To examine the performance of inference with various sizes of models, we conduct experiments under different settings. One model that we use is an 1.3B early-exit model pre-trained from scratch with 300B tokens, and the other one is a 7B model pre-trained with 150B tokens. Each of them has two minimalistic early exits added to the 1/4 and 1/2 depth respectively; other configurations follow those in Section 7.1.1. Neither model has gone through any fine-tuning or alignment. Since designing the decoding method or early-exit condition is not the focus of our work, we simply adopt greedy decoding and a basic confidence-based decision rule for early exiting: at each early exit, if the maximum probability of next-token prediction is above a pre-defined threshold, then the model chooses the most likely token and moves on to generate the next one. The trade-off between output quality and speed of generation is then controlled by the threshold. Note that for a fair comparison, when the threshold is set to 1, early-exit layers are disabled, so that the computational complexity of a full-model forward pass matches that of a standard LLM without early exits; the inference latency with the threshold set to 1 is also used as the baseline for calculating (relative) speedup.
@@ 7.2.1 Output quality and speed of generation
We evaluate the downstream performance of our models with HELM [43] in four tasks. Among them, BoolQ [12] and TruthfulQA [44] are question-answering tasks, while XSUM [46] and CNN/DailyMail [64] are summarization tasks. We use EM (ratio of exact match with the correct answers) as the metric for the former, and ROUGE-L (a measure of similarity to the reference summaries) for the latter. Figure 9 demonstrates the relation between scores and inference speedup for pipeline-based inference with varying thresholds. Encouragingly, for many of these tasks, inference with early exiting can achieve 2×2$\times$2 × or higher speedup compared to full-model inference, with comparable or even better evaluation scores. Table 2 provides a concrete piece of text generated by our 7B early-exit model, showing that speedup can be achieved with minor or no impact on the resulted sequence. Table 3 goes one step further and demonstrates the confidence at each exit for each token, confirming the existence of easy tokens that can be predicted with high confidence at early exits, without wasting computational resources on later layers. It is worth noting that performance of early-exit inference might be improved in many ways, e.g. using early exits of different structures, other decoding mechanisms and exit conditions, or simply more extensive and sufficient pre-training. We also conjecture that early exiting can, in principle, bring higher speedup for larger models, as expressivity of early exits grows with the model capacity, which allows them to make confident and accurate predictions for a larger proportion of tokens.
@@
Figure 9: Evaluation scores and relative speedup with our early-exit models. For each curve, the confidence threshold decreases from left to right.
Table 2: Sequences generated by our 7B early-exit model for the same prompt, with varying confidence thresholds and inference latencies. Differences from the result of full-model inference are highlighted in red.
Prompt
Artificial General Intelligence is
Full model
(Time: 1.42 s)
a branch of computer science that studies the development of computer systems that can perform tasks that normally require human intelligence.
The term "artificial intelligence" was coined by John McCarthy in 1956.
$\text{Threshold}=0.8$
(Time: 1.17 s)
a branch of computer science that studies the development of computer systems that can perform tasks that normally require human intelligence.
The term "artificial intelligence" was coined by John McCarthy in 1956.
$\text{Threshold}=0.4$
(Time: 0.89 s)
a branch of computer science that studies the development of computer systems that can perform tasks that normally require human intelligence.
The term artificial general intelligence was coined by John McCarthy in 1965.
$\text{Threshold}=0.2$
(Time: 0.58 s)
a branch of computer science that deals with the development of computer programs that can be used to perform tasks such as problem solving, decision making, and learning.
Artificial General Intelligence is a branch of computer science that
[…]
Layer 8
Layer 16
Final layer
located
Be
Be
ij (0.992) ij (0.995) ij (0.996)
ing (0.999) ing (0.999) ing (0.999)
[…]
The
The
The
capital
capital
capital
of
of (0.832) of (0.860)
China
the
the
country
United
United
States (0.880) States (0.816) States (0.858)
is
is
is
Washington (0.877) Washington (0.992) Washington (0.989)
D
D
D
. (0.999) . (0.997) . (0.997)
C (0.969) C (0.967) C (0.934)
. (0.830) . (0.873) . (0.904)
@@ 7.2.2 A comparison between two inference methods
We compare the inference time of both pipeline-based method and KV recomputation. Since the pipeline-based method uses 4 GPUs for a pipeline parallelism (PP) degree of 4, we allow KV recomputation to use a tensor parallelism (TP) degree of 1 or 4 for a fair comparison. The empirical results on XSUM and CNN/DailyMail tasks are shown in Figure 10. We first notice that the pipeline-based approach outperforms KV recomputation with $\text{TP}=1$ for confidence thresholds smaller than 1 (i.e. when early exiting actually happens), despite the point-to-point communication overhead. Moreover, KV recomputation with $\text{TP}=4$ is faster than the pipeline-based approach; it is worth noting, however, that the small speedup from $\text{TP}=1$ to $\text{TP}=4$ for KV recomputation is only possible with high-end hardware like A100 GPUs connected by high-bandwidth communication. In many use cases of LLM inference, such hardware is not available, and pipeline parallelism is the only option for partitioning a model that is too large to fit into the memory of one single device. In sum, the proposed pipeline-based approach is a more practical option for LLM inference with model partitioning in a wide range of scenarios.
@@
8 Related works
As introduced in Section 1, early exiting has been widely applied for accelerating inference of deep neural networks in the literature. In terms of early-exit Transformers in the NLP domain, the majority of prior works is focused on BERT [16] or other encoder-only models for classification tasks [45, 25, 94, 19, 63, 81, 62, 82, 42, 26]. Recent works have begun to study token-wise early exiting for accelerating inference of encoder-decoder or decoder-only LLMs in autoregressive sequence generation [62, 13, 3, 76]. While they are largely focused on designing inference mechanisms, the lack of support in prior works for training early-exit models with massive 3D parallelism inevitably posts an obstacle towards truly scaling up early-exit LLMs. Our work on EE-LLM is one important step towards eliminating this obstacle, and we also contribute a novel pipeline-based inference mechanism along the way.
Interestingly, early exiting has been used for many other purposes in the literature, besides accelerating inference. For example, early-exit losses can provide deep supervision and regularization for training deep networks with enhanced stability of convergence [69, 39], which motivates our implementation of the cool-down option for non-constant early-exit loss weights, as explained in Section 4.3.1. Other works use early exits for interpreting the intermediate features learned by each layer of a Transformer [37], for improving outputs of the final exit [22], or as the draft models in speculative decoding [33]. Early exiting has also found application in long-tailed classification [18], federated learning with heterogeneous clients [29, 93], token reduction [1], layer-wise training of deep networks [60, Section 5.3], among others. We believe that our progress in this work can be beneficial for these various purposes.
It is worth mentioning that there are other types of dynamic neural networks [24, 83] that facilitate conditional computation and elastic inference, such as layer skipping [6, 5, 79, 13, 17, 87, 78] and mixtures of experts [30, 66, 21]. Another line of work aims to accelerate LLM inference by designing new decoding methods rather than model architectures [40, 59, 91]. Each of these approaches has its own pros and cons, and some of them can be complemented by early exiting. A detailed comparison between these methods is beyond the scope of the current work.
@@ 9 Conclusions
We have introduced EE-LLM, a system for large-scale training and inference of early-exit LLMs with 3D parallelism. For training, we have presented how to execute backpropagation of the early-exit training objective across pipeline stages, various performance optimizations for minimizing the computational overhead compared to training standard LLMs, and some advanced features for more fine-grained control and optimization of the training process. For inference, we have introduced our design and implementation of two inference methods, one based on KV recomputation and the other based on a new type of pipeline parallelism, both of which are compatible with KV caching for autoregressive generation. Along the way, we have discovered some interesting chemical reaction between early exiting and pipeline parallelism, which is likely because it is both along the depth dimension that an early-exit model executes forward computation selectively and gets partitioned by pipeline parallelism. With EE-LLM, it now becomes possible to train and deploy early-exit LLMs that are as large as any standard LLM allowed by Megatron-LM, given the same amount of computational resources. We hope that EE-LLM will be a helpful tool for future study and applications of early-exit LLMs at larger scales.
@@
Appendix A Analysis of training efficiency
In this section, we derive formulas for the training time per iteration and peak GPU memory usage across pipeline stages under general configurations of early exits, in terms of the quantities listed in Table 4. Some of these quantities, especially $\{f_{\mathsf{o}},b_{\mathsf{o}},m_{\mathsf{o}},m^{\dagger}_{\mathsf{o}}\}$, can be further calculated analytically with lower-level quantities such as sequence length, microbatch size, hidden size, vocabulary size and others [57, 49, 36], or estimated empirically by profiling in practice.
It is assumed that all pipeline stages have the same number of Transformer layers on the backbone, and all early-exit layers follow the same structure. We also assume that input and output embedding matrices are untied, although it is not hard to extend our analysis to more general cases. For simplicity, we ignore the point-to-point communication latency between pipeline stages in our analysis, since it typically takes a minor proportion of the overall training time, and can often be overlapped with computation.
Table 4: A list of notation for theoretical analysis. The following abbreviations are used: INinput processing layer; EEearly-exit layer; FEfinal-exit layer; BB — Transformer backbone on one pipeline stage.
Notation Definition
$P$ Number of pipeline stages
$N_{i}$ Number of early exits in Stage $i\in[P]$
$M$ Number of microbatches for one training iteration
$f_{\mathsf{o}},b_{\mathsf{o}}$ Forward and backward time of one $\mathsf{o}\in\{\mathsf{IN},\mathsf{EE},\mathsf{FE},\mathsf{BB}\}$ for one microbatch
$m_{\mathsf{o}}$ Memory usage for storing the model parameters of one $\mathsf{o}\in\{\mathsf{IN},\mathsf{EE},\mathsf{FE},\mathsf{BB}\}$
$m^{\dagger}_{\mathsf{o}}$ Activation memory of one $\mathsf{o}\in\{\mathsf{IN},\mathsf{EE},\mathsf{FE},\mathsf{BB}\}$ for one microbatch
$\mathbbm{1}(\cdot)$ The indicator function, $\mathbbm{1}(\mathcal{E})=1$ if the event $\mathcal{E}$ holds true, and 00 otherwise
A.1 Training time per iteration
Let us analyze the training time per iteration step by step.
Step 1: simplified analysis without early exits.
We start by recalling the simplified analysis in prior works for the standard 1F1B pipeline schedule [49], which assumes that all stages have the same forward time $f$ and backward time $b$ for one microbatch. Recall from Figure 3 that the critical path of one training iteration consists of three parts:
  • Part 1: forward steps of the first microbatch on all stages except the last one, which takes time $(P-1)\times f$;
  • Part 2: the steady 1F1B phase with all $M$ microbatches on the last stage, which takes time $M\times(f+b)$;
  • Part 3: backward steps of the last microbatch on all stages except the last one, which takes time $(P-1)\times b$.
Taking the sum of these three parts, we have
time per $\text{time per iteration}=(P-1)\times f+M\times(f+b)+(P-1)\times b=\underbrace{(P-1)\times(f+b)}_{\text{Parts 1 and 3}}+\underbrace{M\times(f+b)}_{\text{Part 2}}.$
Step 2: fine-grained analysis without early exits.
We provide a more fine-grained analysis, by considering the computation related to the input processing layer ($\mathsf{IN}$) and final-exit layer ($\mathsf{FE}$) separately from the Transformer backbone ($\mathsf{BB}$). It is reasonable to assume that $f_{\mathsf{IN}}<f_{\mathsf{FE}}$ and $b_{\mathsf{IN}}<b_{\mathsf{FE}}$, which implies that the last stage is the bottleneck of forward and backward time. In this setting, Parts 1 and 3 of the critical path takes time $f_{\mathsf{IN}}+(P-1)\times f_{\mathsf{BB}}$ and $b_{\mathsf{IN}}+(P-1)\times b_{\mathsf{BB}}$, respectively. Similarly, Part 2 now takes time $M\times(f_{\mathsf{BB}}+b_{\mathsf{BB}}+f_{\mathsf{FE}}+b_{\mathsf{FE}})$. Taking the sum, we arrive at
time per $\text{time per iteration}=\underbrace{f_{\mathsf{IN}}+b_{\mathsf{IN}}+(P-1)\times(f_{\mathsf{BB}}+b_{\mathsf{BB}})}_{\text{Parts 1 and 3}}+\underbrace{M\times(f_{\mathsf{BB}}+b_{\mathsf{BB}}+f_{\mathsf{FE}}+b_{\mathsf{FE}})}_{\text{Part 2}}.$
Step 3: fine-grained analysis with early exits.
Finally, we are ready for our analysis in the setting with early exits ($\mathsf{EE}$). First, it can be checked that early-exit layers incur a total overhead of
${\sum_{i\in[P-1]}N_{i}\times(f_{\mathsf{EE}}+b_{\mathsf{EE}})}$
to Parts 1 and 3. In addition, the sum of forward and backward time of one microbatch for Stage $i$ becomes
$f_{\mathsf{BB}}+b_{\mathsf{BB}}+\mathbbm{1}(i=1)\times(f_{\mathsf{IN}}+b_{\mathsf{IN}})+\mathbbm{1}(i=P)\times(f_{\mathsf{FE}}+b_{\mathsf{FE}})+{N_{i}\times(f_{\mathsf{EE}}+b_{\mathsf{EE}})}.$
Note that Part 2, namely the steady 1F1B phase on the last stage, is now bottlenecked by the maximum forward and backward time across all stages. Putting things together, we have the following upper bound:
time per iteration
$\quad\leq f_{\mathsf{IN}}+b_{\mathsf{IN}}+(P-1)\times(f_{\mathsf{BB}}+b_{\mathsf{BB}})+{\color[rgb]{1,0,0}\sum_{i\in[P-1]}N_{i}\times(f_{\mathsf{EE}}+b_{\mathsf{EE}})}$
$\quad\quad+M\times{\max_{i\in[P]}}\Big{\{}f_{\mathsf{BB}}+b_{\mathsf{BB}}+\mathbbm{1}(i=1)\times(f_{\mathsf{IN}}+b_{\mathsf{IN}})+\mathbbm{1}(i=P)\times(f_{\mathsf{FE}}+b_{\mathsf{FE}})+{\color[rgb]{1,0,0}N_{i}\times(f_{\mathsf{EE}}+b_{\mathsf{EE}})}\Big{\}}.$
The overhead caused by early-exit layers has been highlighted. The first term corresponds to Parts 1 and 3 of the critical path, and the second term corresponds to Part 2. It is worth mentioning that the second term of the overhead might not take effect due to the maximum operator, in which case the early-exit overhead to one training iteration is independent of the number of microbatches.
A.2 Peak GPU memory
Consider the GPU memory for one specific pipeline stage, say Stage $i$. Recall that GPU memory usage during training is mainly composed of memory for model parameters, gradients, optimizer states, and intermediate activations [57, 36]. Among them, the memory for gradients and optimizer states (for many common optimizers, such as SGD with momentum or Adam) is proportional to the number of model parameters. Therefore, we assume that
[…]
for some universal constant $\alpha>1$, whose concrete value depends on the choice of optimizer and numerical precisions. Under this assumption, we have
total $\text{total memory}\approx\alpha\times\text{memory}(\text{model parameters})+\text{memory}(\text{activations}).$ (5)
Now it boils down to deriving the memory for model parameters and activations. Model parameters in Stage $i\in[P]$ include the Transformer backbone ($\mathsf{BB}$), the input processing layer ($\mathsf{IN}$) if $i=1$, the final-exit layer ($\mathsf{FE}$) if $i=P$, and the early exits ($\mathsf{EE}$). In other words, we have
memory(model $\text{memory}(\text{model parameters})=m_{\mathsf{BB}}+\mathbbm{1}(i=1)\times m_{\mathsf{IN}}+\mathbbm{1}(i=P)\times m_{\mathsf{FE}}+{\color[rgb]{1,0,0}N_{i}\times m_{\mathsf{EE}}}.$ (6)
The memory for intermediate activations can be similarly divided into components corresponding to $\mathsf{BB}$, $\mathsf{IN}$, $\mathsf{FE}$ and $\mathsf{EE}$. Note that according to Section 4.2.2, the peak memory usage by activations within one early-exit layer is only $m^{\dagger}_{\mathsf{EE}}$ regardless of the number of in-flight microbatches $P+1-i$. Thus one has
$\text{memory}(\text{activations})=(P+1-i)\times m^{\dagger}_{\mathsf{BB}}+\mathbbm{1}(i=1)\times P\times m^{\dagger}_{\mathsf{IN}}+\mathbbm{1}(i=P)\times m^{\dagger}_{\mathsf{FE}}+{\color[rgb]{1,0,0}N_{i}\times m^{\dagger}_{\mathsf{EE}}}.$ (7)
Combining Eq. (6) and (7) with Eq. (5), we arrive at an estimate of memory usage for Stage $i$:
$\text{total memory}\approx\alpha\times$ $\Big{(}m_{\mathsf{BB}}+\mathbbm{1}(i=1)\times m_{\mathsf{IN}}+\mathbbm{1}(i=P)\times m_{\mathsf{FE}}+{\color[rgb]{1,0,0}N_{i}\times m_{\mathsf{EE}}}\Big{)}$
$+$ $\Big{(}(P+1-i)\times m^{\dagger}_{\mathsf{BB}}+\mathbbm{1}(i=1)\times P\times m^{\dagger}_{\mathsf{IN}}+\mathbbm{1}(i=P)\times m^{\dagger}_{\mathsf{FE}}+{\color[rgb]{1,0,0}N_{i}\times m^{\dagger}_{\mathsf{EE}}}\Big{)}.$
The memory overhead caused by early-exit layers has been highlighted in red. Finally, taking the maximum over $i\in[P]$ gives the peak GPU memory across all pipeline stages.
Appendix B Supplementary materials for Section 4.3.2
@@ B.1 Methodology
The maximum number of microbatches that can be inserted into Part 1 or 2 of the explicit bubbles, without increasing training time per iteration, is $\lfloor(p-1)b/(f+b)\rfloor=\lfloor(p-1)/(f/b+1)\rfloor$, where $f/b$ is (an estimate of) the ratio between forward and backward time. To see this for Part 1 (resp. 2), one simply need to notice that the first (resp. last) pipeline stage has an bubble size $b(p-1)$, while the total forward and backward time for each microbatch is $f+b$. Dividing the first value by the second one concludes the proof.
@@
With $K$ inserted microbatches, the $i$-th microbatch is supposed to go through the forward pass of the first $K+1-i$ stages. However, if there is no early exit on Stage $K+1-i$, then this microbatch only need to go through the forward pass up to the last stage (among the first $K+1-i$ stages) that has at least one early exit.
@@
We can calculate the maximum number of backward stages for each inserted microbatch, while ensuring that no overhead to training time per iteration occurs. Notice that for the $i$-th microbatch, the remaining bubble size after its forward computation at the last stage is $(p-1)b-f-(i-1)(f+b)=pb-i(f+b)$. Dividing this value by $b$ gives the number of backward stages $\lfloor(pb-i(f+b))/b\rfloor=\lfloor p-i(f/b+1)\rfloor$.
@@
The actual implementation of this method in EE-LLM takes a different perspective. Instead of inserting additional microbatches into the original schedule, we keep the number of microbatches per training iteration unchanged, and do the following: (1) we replace the full forward/backward computation of a few microbatches with partial forward/backward computation, which can be placed in the bubble between the warm-up and steady phases; (2) we truncate the backward computation of the last few microbatches. These two steps correspond exactly to Parts 1 and 2 introduced earlier, and the visualization in Figure 5(b) remains valid. Such an implementation reduces the training time per iteration, at the cost of lower data utilization.
B.2 Analysis